OpenStudy (anonymous):

Samuel bought a jacket for $15 in January. The cost of the jacket increased to $27 in September. What percent of the original price is the new price?

OpenStudy (anonymous):

what did u get?

OpenStudy (anonymous):

mmmm thought u had tried

OpenStudy (anonymous):

thats not what i got or any of the suggestions

OpenStudy (anonymous):

x/100*15=27 x=27/15 =1.8 1.8*100=180%

OpenStudy (anonymous):

The number of students who attended Albany Medical College last year was 1200. This year, attendance has increased to 1560 students. What percent of last year’s attendance is this year’s attendance? whats the answer and thank u alot

OpenStudy (anonymous):

130%